The speed of the current in the RAnge River is 3 mph. Marie can kayak 4 miles upstream in the same amount of time it takes her to kayak 10 miles downstream. What is the speek of the kayak if traveling in calm water?
Answer by checkley77(12569)

D=RT
4=(S-3)T OR T=4/(S-3)
10=(S+3)T OR T=10/(S+3)
4/(S-3)=10/(S-3) CROSS MULTIPLY
10(S-3)=4(S+3)
10S-30=4S+12
10S-4S=12+30
6S=42
S=42/6
S=7 MPH FOR THE CANOE IN CALM WATER.
PROOF:
4/(7-3)=10/(7+3)
4/4=10/10
1=1